((x^2)y)+4=(x(y^2))-2, x=? and y=?
I do not know where to start with this type of equation.

Answer by Alan3354(69443)

((x^2)y)+4=(x(y^2))-2, x=? and y=?
I do not know where to start with this type of equation.
------------
You have parentheses that don't contribute anything.
x^2y + 4 = xy^2 - 2, x=? and y=?
-------------
It's 1 equation with 2 variables.
What is it you want to do?
There's no unique solution of x & y.
You can pick any value for x and find y, or v.v.